# Ensemble Watermarks for Large Language Models

### v1

@@ Abstract
The rapid advancement of large language models (LLMs) has made it increasingly difficult to distinguish between text written by humans and machines. While watermarks already exist for LLMs, they often lack flexibility, and struggle with attacks such as paraphrasing. To address these issues, we propose a multi-feature method for generating watermarks that combines multiple distinct watermark features into an ensemble watermark. Concretely, we combine acrostica and sensorimotor norms with the established red-green watermark to achieve a 98% detection rate. After a paraphrasing attack the performance remains high with 95% detection rate. The red-green feature alone as baseline achieves a detection rate of 49%. The evaluation of all feature combinations reveals that the ensemble of all three consistently has the highest detection rate across several LLMs and watermark strength settings. Due to the flexibility of combining features in the ensemble, various requirements and trade-offs can be addressed. Additionally, for all ensemble configurations the same detection function can be used without adaptations. This method is particularly of interest to facilitate accountability and prevent societal harm.
Keywords Watermark  $\cdot$ Stylometry  $\cdot$ Ensemble  $\cdot$ Accountability  $\cdot$ LLM
@@ 1 Introduction
Large language models (LLMs) made major improvements with the inception of encoder and decoder models using attention in the form of transformers (Vaswani et al., 2017). Further refinements were made to LLMs using model pre-training (Devlin et al., 2019) and very large data sets (Radford et al., 2019). Through these developments, it has continuously become harder for humans to differentiate texts written by LLMs from those written by humans, with machine-generated texts sometimes even fooling humans more often than human-written texts (Zellers et al., 2019). While machine detection techniques exist, they are slowly being outgrown by the progress of LLMs, for example, several detection methods for GPT-2 are already shown to struggle with GPT-3 (Fagni et al., 2021). This trend already causes worries in many aspects that will only increase with the trend of improvement and consistent scaling of LLMs (Kaplan et al., 2020). Several potential cases of misuse already exist for these advanced models (Ray, 2023).
@@
Stylometry is the study of assessing a person’s writing style, with many stylometric features being proposed in literature Neal et al. (2017). These features are like a writer’s fingerprint, including syntax, the vocabulary used, the structure and size of sentences, and other idiosyncrasies of the author. All these features can be detected statistically and used in tasks like authorship attribution (Stamatatos, 2009a). However, classic authorship attribution has been shown to struggle with smaller LLMs, such as GPT-2 Uchendu et al. (2020).
Sensorimotor norms are categories based on human cognition. Perceptual modalities like "hearing" or action effectors like "hand" are well researched in psychology and cognitive semantics Lynott et al. (2020), but there is yet little research in computer science. In our approach, a secret key derived from the generated output selects the sensorimotor category and thus influences the generated words, e.g., for the olfactory category "smells funny" would be preferred over "looks funny".
An acrostic is a text in which the first letter of each sentence can be combined to spell out a hidden message or word. Historically, acrostica have been used by authors by encoding their authorship Johnson (2006) with variations of their names being the hidden acrostic. A famous example of an acrostic is reported in Appendix 3. In our approach, the secret key controls the letters to be used as the first letter of the first word of a generated sentence. The main contributions presented in this paper are:
@@
The flexible nature of our ensemble watermark allows it to be adapted to different requirements while using the same detection method.
@@ 2 Background & Related Work
Table 1: Overview of common stylometric features as identified in literature from Niess and Kern (2024). Their suitability for four distinct ways to integrate watermarks into LLMs are reported based on a self-assessment. A black circle shows the best compatibility, followed by the semi-filled circle. For our ensemble we decided on sensorimotoric words and acrostics next to the established red-green feature, and on logit manipulation for watermarking.
@@
Several notable attempts exist to detect machine generated text, albeit without incorporating language and stylometric features. Depending on their functional approach, these efforts can generally be categorized into two primary groups.
@@ 2.1 Watermarking Approaches
In this case, it is assumed that the LLM has already completed text generation sometime in the past. The methods in this group attempt to detect the generated text without employing any watermark or other modifications to either the LLM or the final output. Notable examples of this group are the classifier by OpenAI Kirchner et al. (2023), GPTZero Tian (2023) or DetectGPT (Mitchell et al., 2023) who uses the probability curvature of text sampled from a LLM. DetectGPT utilizes the property that text tends to occupy negative curvature regions of the log probability of the model. This implies that any minor modifications to a sentence will result in a reduction in its log likelihood. This is because an LLM continuously strives to attain the optimal probability for each sentence. More post-hoc methods can be found in the survey by Jawahar et al. (2020). Post-hoc detection methods have the benefit that they can be used on any suspected texts without any prerequisites, but they are susceptible to break by user attacks and become less effective with complex language models Chakraborty et al. (2023)
@@
Watermarking is the technique of hiding information in data that is difficult for others to remove but can be detected by an algorithm to read the hidden information. Some successful watermark implementations already exist, such as Kirchenbauer et al. (2023), Christ et al. (2023), and there is ongoing research by Aaronson (2022, 2023). Specifically, Kirchenbauer et al. (2023), which we also include in our ensemble watermark and all our comparisons, separates tokens into a green and red list, where tokens in the green list get a weight boost to increase their representation in the output. This leads to generated text mostly consisting of words from the green list, while human written text naturally also uses words from the red list. Xiang et al. (2024) address the issue of replacing sensitive words and make use of Sentence-BERT (Reimers and Gurevych, 2019). Zhao et al. (2023) also use red-green watermarking, but instead of dynamically generating the red-green list, use a fixed list to simplify the grouping strategy. Kuditipudi et al. (2024) then test several sampling schemes to improve watermark performance. The Duwak approach Zhu et al. (2024) combines the red-green watermark feature with a sampling scheme to improve watermarking; it does, however, not combine multiple logit watermark features as we do.
@@ 2.2 Stylometric Features
There is a broad range of stylometric features previously proposed in literature (Lagutina et al., 2019; Stamatatos, 2009b). They can broadly be categorized into five levels: lexical, syntactic, semantic, structural, and domain-specific features. In Table 1 we collected in previous work Niess and Kern (2024) popular stylometric features reported in literature, as well as additional watermark features that are considered promising.
@@
Winter (2019) define sensory linguistics as the study of how senses and languages are related. The Lancaster Sensorimotor Norms (Lynott et al., 2020) describes a set of 40,000 sensorimotor words together with a crowed-based assessment into 11 dimensions, forming the base of our sensorimotor features. Words can be represented by six perceptual groups (touch, hearing, smell, taste, vision, interoception) and five actions (mouth/throat, hand/arm, foot/leg, head excluding mouth/throat, and torso). Khalid and Srinivasan (2022) use the Lancaster Norms and find that sensorial language might be intentionally used and cannot be considered a random phenomenon. Perceptional features have also been proposed in literature for authorship attribution, specifically in a cross-language setting (Bogdanova and Lazaridou, 2014).
@@
Stein et al. (2014) considered the task of generating a paraphrased version of an existing text, where the generated text should contain an acrostic. Thereby, they formulated this as a search problem. Shen et al. (2019) utilized a sequence-to-sequence network to generate text incorporating acrostica for English and Chinese. More recently, steganography has been considered to embed secret messages in text using BERT (Yi et al., 2022).
@@ 2.3 Watermarking Implementations
LLMs can utilize watermarks, including stylometric ones, in several ways. We present four different approaches in the following. They have different strengths and weaknesses. We provide an overview and estimate how effective each approach is expected to be for different stylometric features in Table 1.
@@
In our approach, we have opted for directly manipulating the logits of the LLM tokens generated. Logits are the raw output values of a machine learning model before applying an activation function like softmax. They represent the unnormalized score or prediction for each token. This manipulation gives considerable control over how often single features are generated, though the difficulty lies in finding the correct logits that produce the desired features. An example of a watermark of this type is the work by Kirchenbauer et al. (2023).
@@
A standard solution is to conduct additional training to change the output of an LLM. This, of course, also works with watermarks, though it is less controlled than other methods. For example, the written material of an author with a known writing style could be used to fine-tune an LLM to produce text similar to that author Li et al. (2023). The problem is that it is difficult to control what is learned and what is not. The model might learn not only the desired watermark features but also overfit to other features, which can lead to a loss of generalization, domain mismatch, and, in general, reduced robustness for outputting text.
@@
Another option is to use the innate capability of LLMs and to engineer a system prompt that tells the model what features are desired Zhou et al. (2023). For example, it is possible to tell contemporary LLMs to use only certain letters in their sentences OpenAI ChatGPT (2024). The challenges here are creating the right prompt, defending against users writing their own prompt, and determining what prompts the language model can comprehend and adhere to. In general, a more powerful model is more suitable for this method.
@@
The last possibility is to post process the text after generation is finished. This is how early attempts were made by Topkara et al. (2005) and Atallah et al. (2001) when LLMs were not yet developed. This was popular for embedding watermarks at the time, for example, for copyright or document integrity. Although the generative power of LLMs has made this less attractive, it can still be useful for some simple features like synonym replacement that do not require the capability to generate new sentences.
@@ 3 Method
Our approach consists of two main parts: 1) the watermark generation process based on manipulation of logits controlled via dynamic keys, and 2) a test procedure for detecting an existing watermark based on statistical tests.
@@ 3.1 Watermark Generation
We chose a relatively large weight boost for acrostica since the beginning of the sentence is more flexible, and the rest of the sentence can easily adapt to this change. We chose a small weight for sensorimotor words so that the model would not be overly biased.
All features have in common that their strength is controlled via a $\delta$ parameter. In the evaluation we study the impact of these parameters on the generated text.
While we opted for fixed values for the weights, they could also be chosen dynamically, depending on how long it has been since a desired feature was chosen or on the distribution of the current weights.
@@ 3.2 Watermark Detection
The probability of detecting a sensorimotor watermark can use class-specific frequencies from the Google N-gram corpus or evenly split probabilities. For each class $c$, the baseline probability $p_{c}$ is calculated. Given $k_{c}$ matches out of $n_{c}$ words for class $c$, the probability is:
@@ 4.2 Results
Figure 1 provides an overview of combinations of features for Llama 3.1 8B, with a lower Z-Score indicating improved detectability. Overall, combinations of watermark features have a higher detection rate than single features. Of the single features, the red-green feature provides the best detectability.
The combination of all features has the best detection rate, with an increase of 17.78% compared to the red-green feature. We also tested for statistical significance in Figure 3. This finding is consistent across all experiments we conducted and reported in Appendix A for all settings.
@@ 5 Conclusion
We presented a novel ensemble watermark method for large language models, particularly generative transformer models. Out of the possible approaches for integrating watermarks in LLMs we opted for manipulating the probability directly when generating the tokens. Following existing work, we generate a key dynamically derived directly from the generated text to control the watermark features. Out of many possible stylometric features, we focused on two features for our evaluation, the acrostic and sensorimotor norms, and combined them with an established method from literature, red-green lists.
@@ Appendix A Appendix
Table 3: Example of an acrostic written by Arnold Schwarzenegger in a letter to the members of the California State Assembly, which has been used in literature for illustration before [Stein et al., 2014].

### v2

@@ Abstract
As large language models (LLMs) reach human-like fluency, reliably distinguishing AI-generated text from human authorship becomes increasingly difficult. While watermarks already exist for LLMs, they often lack flexibility and struggle with attacks such as paraphrasing. To address these issues, we propose a multi-feature method for generating watermarks that combines multiple distinct watermark features into an ensemble watermark. Concretely, we combine acrostica and sensorimotor norms with the established red-green watermark to achieve a 98% detection rate. After a paraphrasing attack, the performance remains high with 95% detection rate. In comparison, the red-green feature alone as a baseline achieves a detection rate of 49% after paraphrasing. The evaluation of all feature combinations reveals that the ensemble of all three consistently has the highest detection rate across several LLMs and watermark strength settings. Due to the flexibility of combining features in the ensemble, various requirements and trade-offs can be addressed. Additionally, the same detection function can be used without adaptations for all ensemble configurations. This method is particularly of interest to facilitate accountability and prevent societal harm.
Ensemble Watermarks for Large Language Models
Georg Niess1, Roman Kern1,2, 1Graz University of Technology, 2Know Center Research GmbH, Correspondence: georg.niess@tugraz.at
@@ 1 Introduction
The inception of transformer-based architectures (Vaswani et al., 2023) combined with large-scale pertaining (Devlin et al., 2019; Radford et al., 2019) has continuously improved the performance of modern large language models (LLMs). Humans increasingly struggle to distinguish between texts written by LLMs and those created by humans, as machine-generated texts sometimes even deceive people more often than human-written ones (Zellers et al., 2019). Although a growing range of post-hoc detectors exists, many detection methods that were once considered reliable for GPT-2 now struggle with GPT-3 and later versions (Fagni et al., 2021). This arms race between generation and detection intensifies as model size and capabilities continue to scale (Kaplan et al., 2020). Several potential cases of misuse are already associated with these advanced models (Ray, 2023).
@@
Stylometry is the statistical analysis of variations in literary style, with numerous stylometric features proposed in literature Neal et al. (2018). These features serve as a writer’s fingerprint and include aspects such as syntax, vocabulary, sentence structure, sentence length, and other unique authorial characteristics. All of these features can be statistically analyzed and utilized in tasks like authorship attribution (Stamatatos, 2009). However, traditional authorship attribution techniques have shown challenges when applied to even smaller language models like GPT-2 Uchendu et al. (2020).
Sensorimotor norms are classifications based on human cognition. Perceptual modalities, such as "hearing", and action effectors, such as "hand", have been extensively researched in psychology and cognitive semantics  Lynott et al. (2020). However, there is still limited research in computer science on this topic. In our approach, a secret key derived from the generated output selects the sensorimotor category, thereby influencing the generated words. For example, "smells funny" would be preferred over "looks funny" for the olfactory category.
An acrostic is a text in which the first letter of each sentence can be combined to spell out a hidden message or word. Historically, authors have used acrostics to encode their authorship Johnson (2006), often incorporating variations of their names into the hidden message. A notable example of an acrostic can be found in Appendix 4. In our method, the secret key determines which letters are used as the first letter of the first word in each generated sentence.
The main contributions of this paper are as follows:
@@
The flexible nature of our ensemble watermark allows it to be adapted to different requirements while using the same detection method. All code and data generated is available on GitHub [^1] https://github.com/CommodoreEU/ensemble-watermark.
@@ 2 Background & Related Work
Table 1: Overview of common stylometric features as identified in literature. Their suitability for four distinct ways to integrate watermarks into LLMs are reported based on a self-assessment. A black circle shows the best compatibility, followed by the semi-filled circle. For our ensemble we decided on sensorimotoric words and acrostics next to the established red-green feature, and on logit manipulation for watermarking. (Extended from Niess and Kern, 2024)
@@
Several notable attempts have been made to detect machine-generated text without using language and stylometric features. Depending on their functional approach, these efforts can generally be categorized into two main groups.
@@ 2.1 Watermarking Approaches
The techniques in this category aim to identify machine-generated text without adding watermarks or altering either the LLM itself or its output. Notable examples of methods in this include the classifier developed by OpenAI Kirchner et al. (2023), GPTZero Tian (2023), and DetectGPT (Mitchell et al., 2023) which utilizes the probability curvature of text sampled from an LLM. DetectGPT relies on the principle that text typically exists within negative curvature regions of a model’s log probability. This means that even minor changes to a sentence will decrease its log likelihood, as an LLM continuously aims to optimize the probability of each sentence. Additional post-hoc methods can be found in the survey by Jawahar et al. (2020). One advantage of post-hoc detection methods is that they can be applied to any suspected text without prior requirements. However, they are vulnerable to user attacks and become less effective with more complex language models (Chakraborty et al., 2023).
@@
Watermarking is the technique of hiding information in data that is difficult for others to remove but can be detected by an algorithm to read the hidden information. Some successful watermark implementations already exist, such as Kirchenbauer et al. (2023), Christ et al. (2023), and there is ongoing research by Aaronson (2022, 2023). The work by Kirchenbauer et al. (2023), which we also include in our ensemble watermark and all our comparisons, divides tokens into two categories: a green list and a red list. Tokens in the green list are given a weight boost to enhance their representation in the generated output. As a result, the generated text predominantly consists of words from the green list, while human-written text naturally incorporates words from the red list. Additionally, Xiang et al. (2024) tackle the challenge of replacing sensitive words by utilizing Sentence-BERT (Reimers and Gurevych, 2019). Zhao et al. (2024) also use red-green watermarking, but instead of dynamically generating the red-green list, use a fixed list to simplify the grouping strategy. Kuditipudi et al. (2024) then test several sampling schemes to improve watermark performance. The Duwak approach Zhu et al. (2024) combines the red-green watermark feature with a sampling scheme to improve watermarking; it does, however, not combine multiple logit watermark features as we do.
@@ 2.2 Stylometric Features
A wide variety of stylometric features have been proposed in the literature (Lagutina et al., 2019; Stamatatos, 2009). These features can be broadly classified into five categories: lexical, syntactic, semantic, structural, and domain-specific features. In Table 1, we have compiled a list of popular stylometric features found in the literature, along with additional watermark features we consider promising.
@@
Winter (2019) define sensory linguistics as the study of the relationship between senses and language. The Lancaster Sensorimotor Norms (Lynott et al., 2020) detail a set of 40,000 sensorimotor words, accompanied by a crowd-based assessment across 11 dimensions, which serve as the foundation for our sensorimotor features. Every word can be categorized into six perceptual groups: touch, hearing, smell, taste, vision, and interoception, as well as five action categories: mouth/throat, hand/arm, foot/leg, head (excluding mouth/throat), and torso. Khalid and Srinivasan (2022) utilize the Lancaster Norms and conclude that sensorial language is likely used intentionally and should not be viewed as a random phenomenon. Additionally, perceptual features have been suggested in the literature for authorship attribution, particularly in cross-language contexts  (Bogdanova and Lazaridou, 2014).
@@
Stein et al. (2014) explored the task of generating paraphrased versions of existing texts that also include acrostics. They approached this challenge as a search problem. Shen et al. (2019) employed a sequence-to-sequence network to generate texts with acrostics in both English and Chinese. More recently, using steganography for embedding secret messages in text has been investigated, utilizing BERT (Yi et al., 2022).
@@ 2.3 Watermarking Implementations
LLMs can utilize watermarks, including stylometric ones, in several ways. We present four different approaches in the following. They have different strengths and weaknesses. We outline four distinct approaches below, each with its own strengths and weaknesses. Additionally, we provide an overview of these methods and assess their expected effectiveness for different stylometric features in Table 1.
@@
In our approach, we have opted for directly manipulating the logits of the LLM tokens generated. Logits are the raw output values of a machine learning model before applying an activation function like softmax. They represent the unnormalized score or prediction for each token. This manipulation gives considerable control over how often single features are generated, though the difficulty lies in finding the correct logits that produce the desired features. An example of this type of watermark can be found in the work by Kirchenbauer et al. (2023).
@@
A standard solution is to conduct additional training to change the output of an LLM. This method can also be applied to watermarks, although it is less controlled than other techniques. For instance, the written works of an author with a distinctive writing style can be used to fine-tune an LLM to produce text resembling that author’s style Li et al. (2023). However, a significant challenge is controlling what the model learns. The LLM might not only pick up the desired watermark features but could also overfit to unrelated aspects, leading to a loss of generalization, domain mismatch, and decreased robustness in generating text.
@@
Another alternative is to use the inherent capabilities of LLMs by designing a system prompt that specifies the desired features Zhou et al. (2023). For example, it is possible to tell contemporary LLMs to use only certain letters in their sentences OpenAI ChatGPT (2024). The primary challenges associated with this method include creating an effective prompt, protecting against users writing their own prompt, and understanding what prompts the language model can comprehend and adhere to. Generally, more powerful models are better suited for this approach.
@@
The final option is to process the text after generation is complete. This is how early attempts were made by Topkara et al. (2005) and Atallah et al. (2001) before LLMs were developed. It was commonly used to embed watermarks for copyright and document integrity purposes. Although the generative capabilities of LLMs have diminished the appeal of this approach, it can still be beneficial for implementing simple features like synonym replacement, which do not require generating new sentences.
@@ 3 Method
Our approach includes two main components: 1) a process for generating watermarks by manipulating logits with dynamic keys, and 2) a test procedure for detecting an existing watermark using statistical tests.
@@ 3.1 Watermark Generation
We decided to assign a relatively large weight boost to acrostica because the beginning of the sentence is more flexible, allowing the rest of the sentence to adapt to this change easily. In contrast, we chose a small weight for sensorimotor words to prevent the model from becoming overly biased.
All features have in common that their strength is controlled via a $\delta$ parameter. In the evaluation we study the impact of these parameters on the generated text. While we opted for fixed values for the weights, they could also be chosen dynamically, depending on how long it has been since a desired feature was chosen or on the distribution of the current weights.
@@ 3.2 Watermark Detection
The probability of detecting a sensorimotor watermark uses class-specific frequencies from the Google N-gram corpus or evenly split probabilities. For each class $c$, the baseline probability $p_{c}$ is calculated. Given $k_{c}$ matches out of $n_{c}$ words for class $c$, the probability is:
@@ 4.2 Results
Figure 1 provides an overview of combinations of features for Llama 3.1 8B, with a lower Z-Score indicating improved detectability. Overall, combinations of watermark features have a higher detection rate than single features. Of the single features, the red-green feature provides the best detectability. The combination of all features has the best detection rate, with an increase of 17.78% compared to the red-green feature. We also tested for statistical significance in Figure 3. This finding is consistent across all experiments we conducted and reported in Appendix A for all settings.
@@ 5 Conclusion
We introduced a novel ensemble watermarking method for large language models, specifically generative transformer models. Among the various approaches for integrating watermarks into LLMs, we chose to manipulate the probabilities directly when generating tokens. Following existing work, we generate a key dynamically derived directly from the generated text to control the watermark features. Out of many possible stylometric features, we focused on two features for our evaluation, the acrostic and sensorimotor norms, and combined them with an established method from literature, red-green lists.
@@
6 Limitations and Risks
Currently, stylometric watermarks are only implemented for the English language. In principle, our watermark can be applied to almost all languages with two restrictions based on the stylometric features we use. Acrostica require the language to have an alphabet similar in size to the Latin alphabet, so Chinese or Japanese characters would make this aspect nonfunctional. As for sensorimotor norms, all languages share this aspect since they relate directly to how our brains work Connell et al. (2012). A different database would be ideal for each language, although the English one by Lynott et al. (2020) could also be translated for the same effect, since sensorimotor accuracy is not important for watermarking purposes.
This also leads to the topic of attacks. While paraphrasing attacks are commonly considered in literature, other attacks are also plausible.
Another possible limitation is as to what language models are compatible to this method. Currently, we have only shown compatibility to the popular decoder models Llama 3.1/3.2 and Mistral. However, the only architecture limitation for a language model to be compatible is to have a logits layer to manipulate generation and for the model to be sequential to be able to generate new keys.
Risks.
Our work contributes to trustworthy AI, in particular by addressing the topic of accountability for proprietary language models. Since the key generation can be made specific for each LLM, this allows to link generated text back to its origin. Thus, the proposed watermark and its methods have high potential to help reduce the negative risks on society by the harmful use of LLMs.
@@ Appendix A Appendix
Configuration p-Senso p-Acro p-Red-Green
Human 0.60140883 0.99999999 0.94382440
Senso 0.00011047 0.99999999 0.88591727
Acro 0.63667415 0.00000931 0.94849433
Red-Green 0.63739815 0.99999999 0.00000102
Senso + Acro 0.00001441 0.00004255 0.91098964
Senso + Red-Green 0.00034610 0.99999999 0.00000011
Acro + Red-Green 0.58650698 0.00022102 0.00000004
All Three 0.00002030 0.00022102 0.00000001
Table 3: Table of p-values of each feature under different ensemble settings. All values are from Llama 8B with the medium strength watermark setting.
Table 4: Example of an acrostic written by Arnold Schwarzenegger in a letter to the members of the California State Assembly, which has been used in literature for illustration before (Stein et al., 2014).